OpenStudy (jamail21334):

It takes Nate five hours to paint a wall, and it takes Jill seven hours to paint the same wall. About how many hours will it will take if Nate and Jill work together to paint the wall?

OpenStudy (whpalmer4):

This is a math question, not a music question, but I'll answer anyhow. One very effective way to work these problems is to look at unit rates. Nate paints a wall in 5 hours, which means in 1 hour, he paints 1/5 of a wall (1 wall / 5 hours = 1/5 wall/hour). Similarly, Jill works at a rate of 1/7 wall/hour. Together, assuming as these problems do that they can work together with no loss of efficiency, they will do \[\frac{1}{5} + \frac{1}{7}\] walls every hour. \[\frac{1}{5} + \frac{1}{7} = \frac{7}{7}*\frac{1}{5} + \frac{5}{5}*\frac{1}{7}\](making a common denominator) \[\frac{7}{7}*\frac{1}{5} + \frac{5}{5}*\frac{1}{7} = \frac{7}{35} + \frac{5}{35} = \frac{7+5}{35} = \frac{12}{35}\] Nate and Jill working together will complete \(\dfrac{12}{35}\) of a wall every hour. You'll need to figure out from that how many hours it will take to do the whole wall, if 1 hour gets you \(\dfrac{12}{35}\) of a wall.